Aman left 1/3rd of his money to his daughter,1/4th to his son , and the remaining 20,000 to his wife. How much money did the man have??
Answer by tommyt3rd(5050)

1/3+1/4=
5/12
so T- 7/12 T =20000
5/12T=20000
T=48000